# Trust-Oriented Adaptive Guardrails for Large Language Models

### v2

@@ 1 Introduction
Large language models (LLMs), notably GPT, Llama and Claude, have catalyzed innovation across many critical sectors including healthcare, finance, and engineering Ullah et al. (2024). However, with these advancements come significant challenges, particularly concerning about the reliability, safety, robustness and ethical use of these models Huang et al. (2024). Prominent issues such as hallucination, biases, and toxicity highlight the critical need for robust safeguards, a.k.a. guardrails [^1] Details of related work are shown in Appendix A., to ensure that LLMs operate within acceptable boundaries, and it has been argued that the design of such guardrails need to be sociotechnical in their design, considering both ‘social’ and ‘technical’ aspects Conitzer et al. (2024); Dong et al. (2024a). To this end, this paper revisits the trust modeling Yamamoto (1990); Cho et al. (2015) in the context of LLM safeguarding, and integrates it into the guardrail design in order to boost the guardrail’s ability to cater for different needs from a diverse set of users.
In addition to the above ‘social’ need, this work also addresses a ‘technical’ concern that existing guardrails are rule-based Rebedea et al. (2023); Inan et al. (2023) or incorporating extensive post-processing checks Yang et al. (2024) in LLMs applications. While these approaches offer some level of control, they are often static, inflexible and unable to adapt to the dynamic nature of human activities. As a result, there is a growing need for more adaptive and context-sensitive solutions that can better manage the risks associated with deploying LLMs in real-world scenarios.
@@
This paper explores trust modeling [^1] Details of related work are shown in Appendix A., a mechanism that is perceived as the basis for decision making in many contexts for maintaining long-term relationships based on cooperation and collaboration Cho et al. (2015), to address these issues. In other safety-critical areas, trust is utilized to adapt systems’ capabilities, ensuring consistent, safe, and efficient decisions when interacting with users Okamura and Yamada (2020); Wang et al. (2018). Nevertheless, directly implementing trust modeling in LLMs’ guardrails faces several challenges: 1) user profiles are normally private for LLM applications; 2) the trust dimension is difficult to quantify across various application contexts; 3) Difficulty exists in accurately assessing trust during dynamic interactions while balancing risk with data accessibility. In addition to trust modeling, In-context Learning (ICL) via Retrieval-Augmented Generation (RAG) [^1] Details of related work are shown in Appendix A. plays a crucial role in enhancing the adaptability of guardrails. It empowers models to swiftly adapt their outputs without updating parameters, guided by specific instructions or contextual examples presented by trust-oriented retrieved context Fan et al. (2024). Leveraging this capability, it is possible to implement guardrails that are not only adaptive but also finely responsive to the sensitive queries encountered by the LLM, guided by hierarchical knowledge sources and aligned with user trust levels. This synergy between trust modeling and ICL offers a novel framework for ensuring the safety of LLMs while maintaining practicality.
@@ 2 Problem Formulation
Existing guardrails, represented as $G$, often enforce uniform and rigid rules to LLMs’ response on $x$ against a set of predefined safety conditions $C$. This indiscriminate approach to safety may lead to overly restrictive systems that implement moderated responses $M(x)$, such as “I am sorry…”, which do not adequately balance safety with usability and flexibility for diverse user groups. Formally,
$G(x)=\begin{cases}x&\text{if }x\models C\\ M(x)&\text{if }x\not\models C\end{cases}$ (1)
@@
$AG(x,T_{i})=\begin{cases}R(x,T_{i})&\text{if }x\models C\text{ or }(x\not\models C\text{ and }T_{i}\geq\beta)\\ M(x)&\text{if }x\not\models C\text{ and }T_{i}<\beta\end{cases}$ (2)
where $\beta$, initially assigned with a unified base value, serves as a trust threshold dynamically adjusted for different user groups and queries. Intuitively, for users verified by reliable authorities, $\beta$ is relaxed based on authority levels, while for less reliable authorities, $\beta$ is elevated according to query sensitivity to prevent misuse. $R(\cdot)$ is the content richness function, defined in Equation (15) to determine the level of sensitive information provided by knowledge base with hierarchal confidentiality. From Equation (2), it is clear that we keep the strict forbidden for users with low trust levels, but release the restriction for trustworthy user groups to ensure access.
@@ 3 Methodology
Inspired by Cho et al.’s work Cho et al. (2015), which emphasizes the use of weighted aggregation to construct multi-dimensional trust in human-machine interactions, we adopt their classification of trust into communication, information, social, and cognitive dimensions to derive a composite trust score. Each dimension includes multiple trust attributes, from which we selected the most applicable ones for the adaptive guardrail context, with the flexibility to integrate others whenever needed. We design two modules following Parhizkar et al. (2021), i.e., Direct Interaction Trust ($DT$) and Authority-Verified Trust ($AT$). $DT$ is crucial in environments where advisors are dishonest or exhibit frequently changing behaviors, as it captures communication and information trust through user interaction history and content assessments. In contrast, $AT$ as indirect trust is more effective in contexts with reliable advisors, as it evaluates social and cognitive trust using third-party verification. By integrating real-time direct trust assessments with robust external validations, these modules are essential for ensuring the system remains adaptive and reliable Jøsang et al. (2007). Eventually, if the composite trust score calculated by these two modules surpasses a predefined threshold, LLMs can adaptively adjust their responses to enhance system integrity; otherwise, access is restricted to prevent potential misuse.
@@
$T_{i}=\eta\cdot AT_{i}+(1-\eta)\cdot DT_{i}$ (3)
where $\eta$ is an adaptive weight that adjusts based on the authority rankings of TTPs and the threshold of direct interaction trust, $\delta$. This mechanism highlights that in environments with reliable TTPs, indirect trust can be more beneficial than direct trust Parhizkar et al. (2021). On the other hand, when the assurance from TTPs is less reliable, the guardrails augment the $DT$ weight to ensure the security and integrity of LLMs’ inputs and responses, effectively managing access to sensitive information during dynamic interaction. The adaptive weight is formulated as below:
@@
3.0.1 Direct Interaction Trust
Based on Slovic (1993); Siegrist (2021); Frisch and Giulianelli (2024), Direct Interaction Trust should satisfy the following properties: 1) Trust gradually decreases over time, exhibiting a temporal correlation that strengthens as it approaches the present moment. 2) Additionally, trust should be strongly correlated with the consistency of interactions, and trust is more readily destroyed by negative experiences. To achieve them, we formulate a mathematical expression in Eq. (9), related to time and consistency during interactions, which is used to quantitatively compute the DT.
@@ 3.0.1 Direct Interaction Trust
where $a^{(\tau_{t})}_{i}$ and $b^{(\tau_{t})}_{i}$ represent the counts of safe and unsafe interactions at each past interaction $t$, within the sliding window $W$ which limits the summation to the most recent $W$ interactions. $safe^{(\tau)}$ and $unsafe^{(\tau)}$ are the counts of safe and unsafe interactions recorded at the current time $\tau$.
Moreover, in interactions involving LLMs, consistency always serves as a critical factor of interaction uniformity. It quantifies the alignment of a current interaction with historical interactions, providing insights into user behavior patterns. We define the interaction consistency score by applying a quadratic transformation to the cosine similarity metric, which compares embedding vectors from current and past interactions Zeng et al. (2024). This transformation effectively scales the values to remain within the $[0,1]$ interval, while enhancing their directional alignment:
$IC^{(\tau)}_{i}=\frac{1}{W}\sum_{t=1}^{W}\left(\frac{1+\frac{\mathbf{v}\cdot\mathbf{v}_{t}}{\|\mathbf{v}\|\|\mathbf{v}_{t}\|}}{2}\right)^{2}$ (8)
where $\mathbf{v}$ is the feature vector representing the current interaction. $\mathbf{v}_{t}$ represents the feature vector of the $t$-th historical interaction within the sliding window.
Finally, based on the above ingredients, we adopt Bayesian inference formula based on Beta probability density function to evaluate the direct interaction trust, inspired by Josang and Ismail (2002); Chen et al. (2021). The interaction history and consistency score are combined to form direct interaction trust to offer a complete perspective for evaluating continuous and dynamic user engagement:
@@
3.0.2 Authority-Verified Trust
Authority-verified trust utilizes demonstrations from Trust Third Parties (TTPs). Inspired by prior research on recommendation systems Chen et al. (2021), we introduce a trust algorithm and reinterpret it to address the unique needs of LLM guardrails, which incorporates similarity of views ($S$) and confidence level ($C$) as shown in Eq. (13). Specifically, $S$ is based on the principle that a trustor is more likely to accept suggestions from TTPs whose perspectives resonate more closely with its own evaluations Fan et al. (2014). In our context, we assess $S$ by comparing the historical average direct interaction trust and the normalized rating credits ($NR$) assigned by TTPs to corresponding trustees.
Additionally, confidence level $C$ is also a significant factor that should be considered in trust dynamics, as trust relies on the stability and reliability of expectations, which are determined by the level of confidence Luhmann and others (2000); Urbano et al. (2009). It can be gauged by the volume of trust attributes that TTPs attribute to trustees, where a higher positive count indicates greater trust. Trust attributes encompass factors such as social reputation, job title, historical credibility, and family background, alongside other pertinent factors like professional achievements and peer evaluations. They present a comprehensive view of a user’s reliability and standing within their social networks. For the relevance and timeliness of these attributes, they are periodically refreshed through updated data requests from TTPs. The accumulated positive ($pos_{pu}$) and negative ($neg_{pu}$) trust attributes from a TTP $p_{k}$ about the $u_{i}$ are represented statistically and responsible for calculating TTPs’ confidence $C$, allowing quantify trust based on the diversity of trust attributes $j$.
$pos_{ki}^{(\tau)}=\mathcal{A}\cdot\sum_{j=1}^{J}pos_{ki}^{j}$ (10)
$neg_{ki}^{(\tau)}=\mathcal{A}\cdot\sum_{j=1}^{J}neg_{ki}^{j}$ (11)
To enhance the precision of area relevance ($AR$) assessments within our trust model, we employ Prompt Engineering to direct the inference process of multiple LLMs debating $\mathcal{D}$. The robustness of our final decision is strengthened by integrating semantic similarity measures, denoted as $ST$. Specifically, we employ the all-mpnet-base-v2 sentence transformer in our cases, which yields semantically rich embeddings that effectively capture nuanced linguistic relations Reimers and Gurevych (2019). This method refines the evaluation of area relevance between a trustee’s professional area, as verified by TTPs, and the sensitivity required by the prompts. The formula for computing $AR$ in $[0,1]$ is presented as follows:
@@ 3.0.2 Authority-Verified Trust
Remark 1: We also incorporate periodic re-validation throughout the whole trust evaluation process, which is a common practice in web services Ceccarelli et al. (2015), to ensure that all sensitive information requests meet specific security requirements within guardrail systems. This mandatory procedure applies universally, irrespective of trust scores, preventing outdated or overly trusted credentials from granting inappropriate access and thus enhancing the system’s robustness. For long-term access to sensitive or controversial information for legal or specialized purposes, users must periodically re-validate their identity with TTPs to maintain credential validity.
@@ 3.1 Contextual Adaptation Mechanism
Guardrails % CS-related Questions % Irrelevant Sensitive Questions
@@ 4.1 Data Preparation
For our experiments, we adapted AdvBench dataset Zou et al. (2023), categorizing prompts by their relevance to computer science (CS) to evaluate our adaptive trust model for guardrails. This dataset comprises 520 harmful prompts reflecting potential sensitive interactions, supplemented by 3,000 safe prompts for interaction diversity. We specifically selected 162 harmful prompts highly relevant to CS and 358 from other domains to test the guardrail’s efficacy in managing access to sensitive information. Due to privacy concerns and restricted access to actual user data from TTPs, we created synthetic profiles for virtual users, incorporating simulated ratings, professional areas, social information, and more, to emulate real-world trust attributes, allowing us to verify if the trust model effectively aligns access to sensitive information with validated user identities. All necessary materials including code and datasets, are hosted on GitHub [^4] The anonymized GitHub will be made public upon paper acceptance to comply with double-blind review requirements.
4.2 Implementation Assumptions

### v3

@@ 1 Introduction
Large language models (LLMs), such as GPT [1] and Llama [37], have significantly expanded the reach of Artificial Intelligence (AI) technologies across critical sectors such as healthcare, finance, and engineering [38]. However, with these advancements come significant challenges, particularly concerning the reliability, safety, robustness, and ethical use of such models [16]. Prominent issues such as hallucination, biases, and toxicity highlight the critical need for robust safeguards, a.k.a. guardrails, to ensure that LLMs operate within acceptable boundaries, and it has been argued that the design of such guardrails need to be sociotechnical in their design, considering both ‘social’ and ‘technical’ aspects [5, 7]. To this end, this paper revisits the trust modeling [41, 4] in the context of LLM safeguarding, and integrates it into the guardrail design in order to boost the guardrail’s ability to cater for different needs from a diverse set of users.
In addition to the above ‘social’ need, this work also addresses a ‘technical’ concern that existing guardrails are rule-based [31, 17] or incorporating extensive post-processing checks [42] in LLMs applications. While these approaches offer some level of control, they are often static, inflexible and unable to adapt to the dynamic nature of human activities. As a result, there is a growing need for more adaptive and context-sensitive solutions that can better manage the risks associated with deploying LLMs in real-world scenarios.
@@
This paper explores trust modeling, a mechanism that is perceived as the basis for decision making in many contexts for maintaining long-term relationships based on cooperation and collaboration [4], to address these issues. In other safety-critical areas, trust is utilized to adapt systems’ capabilities, ensuring consistent, safe, and efficient decisions when interacting with users [28, 40]. Nevertheless, directly implementing trust modeling in LLMs’ guardrails faces several challenges: 1) user profiles are normally private for LLM applications; 2) the trust dimension is difficult to quantify across various application contexts; 3) Difficulty exists in accurately assessing trust during dynamic interactions while balancing risk with data accessibility. In addition to trust modeling, In-context Learning (ICL) via Retrieval-Augmented Generation (RAG) plays a crucial role in enhancing the adaptability of guardrails. It empowers models to swiftly adapt their outputs without updating parameters, guided by specific instructions or contextual examples presented by trust-oriented retrieved context [11]. Leveraging this capability, it is possible to implement guardrails that are not only adaptive but also finely responsive to the sensitive queries encountered by the LLM, guided by hierarchical knowledge sources and aligned with user trust levels. This synergy between trust modeling and ICL offers a novel framework for ensuring the safety of LLMs while maintaining practicality.
@@
2 Related Work
2.1 Existing Guardrails
Guardrails for LLMs are programmable, rule-based systems acting as intermediaries to ensure that interactions between users and models adhere to established ethical and operational standards [8]. Reflecting the latest advancement, several existing techniques represent the state of the art in this field. Llama Guard [17] operates as a binary classifier that effectively differentiates between safe and unsafe content, ensuring moderation aligns with predefined safety guidelines. Concurrently, Nvidia NeMo [31] enhances interaction accuracy by utilizing the K-nearest neighbors method to align user intents with the most appropriate vector-based canonical forms. After this matching process, LLMs generate safe and pertinent responses, guided by Colang scripts. Moreover, Guardrails AI [30] bolsters system reliability and trust by implementing structured safety protocols, involving detailed XML specifications that rigorously check that outputs adhere to strict content integrity and compliance standards. Supporting these technologies, Python packages such as LangChain [^3] https://www.langchain.com/, AI Fairness 360444https://github.com/Trusted-AI/AIF360, Adversarial Robustness Toolbox [^5] https://github.com/Trusted-AI/adversarial-robustness-toolbox, Fairlearn [^6] https://github.com/fairlearn/fairlearn, and Detoxify [^7] https://github.com/unitaryai/detoxify address biases, boost robustness, and ensure content safety and fairness, collectively fostering responsible and reliable LLM applications.
2.2 Trust Models
Trust is defined by the Merriam-Webster Dictionary [6] as “assured reliance on the character, ability, strength, or truth of someone or something," which succinctly encapsulates trust as a fundamental aspect of relationships, where a trustor places confidence in a trustee based on established criteria. Recognized as a pivotal element in decision-making, trust spans various disciplines including management, psychology, economics, and engineering, underscoring its multidisciplinary importance [21].
Recently, trust in human-technology relationships has gained widespread attention and is typically classified into three distinct types: credentials-based, cognitive trust, and experience-based [15]. Credentials-based trust is commonly employed in security systems, where users are required to provide valid credentials to verify their identity and gain access to services, satisfying established security policies [25]. Cognitive trust involves in the human aspects of trust, focusing on subjective judgments and trust-based decisions in interactions, especially between humans and robots, highlighting the psychological dimensions of trust [9]. Experience-based trust is frequently used in multiagent systems and e-service platforms, where a trustee’s reliability is assessed based on past interactions and reputation-based trust, helping to predict the trustees’ future behaviors [13, 33]. This approach helps predict future behaviors by employing statistical methods to calculate trust scores. In this paper, we introduce a composite trust model that identifies relevant dimensions from various trust paradigms to meet the dual needs of security and practicality in LLM-user interactions, comprehensively assessing trustee reliability in LLMs equipped with adaptive guardrails.
2.3 In-Context Learning
In-context learning utilizes the generalization capabilities of LLMs to efficiently perform tasks with just a few contextual examples, requiring minimal data and no parameter updates, setting it apart from standard in-weights learning like gradient-based fine-tuning [20]. This method has become a focal point due to its ability to streamline the learning process and reduce the computational overhead typically associated with training large models. Research has focused on the selection of demonstration examples and the better ways of problem formulations [44, 23]. A common enhancement to ICL is RAG, which retrieves external knowledge to refine LLM outputs with contextually rich information, improving accuracy and adaptability in real-world scenarios [35, 11]. In our paper, we extend this approach by leveraging RAG to implement online ICL, dynamically retrieving trust-oriented information from hierarchical knowledge bases. By incorporating trust-based criteria, our method ensures that the retrieved knowledge is not only contextually rich but also aligned with user-specific constraints, enabling LLMs to interact effectively with external agents or access expert-curated knowledge sources.
@@ 3 Problem Formulation
Existing guardrails, formally represented as $G$, often enforce uniform and rigid rules to LLMs’ response on $x$ against a set of predefined safety conditions $C$. This indiscriminate approach to safety may lead to overly restrictive systems that implement moderated responses $M(x)$, such as “I am sorry…", which do not adequately balance safety with usability and flexibility for diverse user groups.
$\small G(x)=\begin{cases}x&\text{if }x\models C\\ M(x)&\text{if }x\not\models C\end{cases}$ (1)
@@
$\small AG(x,T_{i})=\begin{cases}R(x,T_{i})&\text{if }x\models C\text{ or }(x\not\models C\text{ and }T_{i}\geq\beta)\\ M(x)&\text{if }x\not\models C\text{ and }T_{i}<\beta\end{cases}$ (2)
where $\beta$, initially assigned with a unified base value, serves as a trust threshold dynamically adjusted for different user groups and queries. Intuitively, for users verified by reliable authorities, $\beta$ is relaxed based on authority levels, while for less reliable authorities, $\beta$ is elevated according to query sensitivity to prevent misuse. $R(\cdot)$ is the content richness function, defined in Equation (14) to determine the level of sensitive information provided by knowledge base with hierarchal confidentiality. From Equation (2), it is clear that we keep the strict forbidden for users with low trust levels, but release the restriction for trustworthy user groups to ensure personalized access.
@@ 4 Methodology
Inspired by Cho et al.’s work [4], which emphasizes the use of weighted aggregation to construct multi-dimensional trust in human-machine interactions, we adopt their classification of trust into communication, information, social, and cognitive dimensions to derive a composite trust score. Each dimension includes multiple trust attributes, from which we selected the most applicable ones for the adaptive guardrail context, with the flexibility to integrate others whenever needed. We design two modules following [29], i.e., Direct Interaction Trust ($DT$) and Authority-Verified Trust ($AT$). $DT$ is crucial in environments where advisors are dishonest or exhibit frequently changing behaviors, as it captures communication and information trust through user interaction history and content assessments. In contrast, $AT$ as indirect trust is more effective in contexts with reliable advisors, as it evaluates social and cognitive trust using third-party verification. By integrating real-time direct trust assessments with robust external validations, these modules are essential for ensuring the system remains adaptive and reliable [19]. Eventually, if the composite trust score calculated by these two modules surpasses a predefined threshold, LLMs can adaptively adjust their responses to enhance system integrity; otherwise, access is restricted to prevent potential misuse.
@@
$\small T_{i}=\eta\cdot AT_{i}+(1-\eta)\cdot DT_{i}$ (3)
where $\eta$ is an adaptive weight that adjusts based on the authority rankings of Trust Third Parties (TTPs) and the threshold of direct interaction trust, $\delta$. This mechanism highlights that in environments with reliable TTPs, indirect trust can be more beneficial than direct trust [29]. On the other hand, when the assurance from TTPs is less reliable, the guardrails augment the $DT$ weight to ensure the security and integrity of LLMs’ inputs and responses, effectively managing access to sensitive information during dynamic interaction. The adaptive weight is formulated as below:
@@
4.1 Direct Interaction Trust
Based on [36, 34, 12], Direct Interaction Trust should satisfy the following properties: 1) Trust gradually decreases over time, exhibiting a temporal correlation that strengthens as it approaches the present moment. 2) Additionally, trust should be strongly correlated with the consistency of interactions, and trust is more readily destroyed by negative experiences. To achieve them, we formulate a mathematical expression in Eq. (9), related to time and consistency during interactions, which is used to quantitatively compute the DT.
@@ 4.1 Direct Interaction Trust
where $a^{(\tau_{t})}_{i}$ and $b^{(\tau_{t})}_{i}$ represent the counts of safe and unsafe interactions at each past interaction $t$ within the sliding window $W$, which limits the summation to the most recent $W$ interactions. $safe^{(\tau)}$ and $unsafe^{(\tau)}$ are the corresponding counts at the current time $\tau$.
Moreover, in interactions involving LLMs, consistency always serves as a critical factor of interaction uniformity. It quantifies the alignment of a current interaction with historical interactions, providing insights into user behavior patterns. We define the interaction consistency score by applying a quadratic transformation to the cosine similarity metric, which compares embedding vectors from current and past interactions [43]. This transformation effectively scales the values to remain within the $[0,1]$ interval, while enhancing their directional alignment:
$\small IC^{(\tau)}_{i}=\frac{1}{W}\sum_{t=1}^{W}\left(\frac{1+\frac{\mathbf{v}\cdot\mathbf{v}_{t}}{\|\mathbf{v}\|\|\mathbf{v}_{t}\|}}{2}\right)^{2}$ (8)
where $\mathbf{v}$ is the feature vector representing the current interaction. $\mathbf{v}_{t}$ represents the feature vector of the $t$-th historical interaction within the sliding window. Finally, based on the above ingredients, we adopt Bayesian inference formula based on Beta probability density function to evaluate the direct interaction trust, inspired by [18, 3]. The interaction history and consistency score are combined to form direct interaction trust to offer a complete perspective for evaluating continuous and dynamic user engagement:
@@
4.2 Authority-Verified Trust
Authority-verified trust utilizes demonstrations from Trust Third Parties (TTPs). Inspired by prior research on recommendation systems [3], we introduce a trust algorithm and reinterpret it to address the unique needs of LLM guardrails, which incorporates similarity of views ($S$) and confidence level ($C$) as shown in Eq. (12). Specifically, $S$ is based on the principle that a trustor is more likely to accept suggestions from TTPs whose perspectives resonate more closely with its own evaluations [10]. In our context, we assess $S$ by comparing the historical average direct interaction trust and the normalized rating credits ($NR$) assigned by TTPs to corresponding trustees.
Additionally, confidence level $C$ is also a significant factor that should be considered in trust dynamics, as trust relies on the stability and reliability of expectations, which are determined by the level of confidence [24, 39]. It can be gauged by the volume of trust attributes that TTPs attribute to trustees, where a higher positive count indicates greater trust. Trust attributes encompass factors such as social reputation, job title, historical credibility, and family background, alongside other pertinent factors like professional achievements and peer evaluations. They present a comprehensive view of a user’s reliability and standing within their social networks. For the relevance and timeliness of these attributes, they are periodically refreshed through updated data requests from TTPs. The accumulated positive ($pos_{pu}$) and negative ($neg_{pu}$) trust attributes from a TTP $p_{k}$ about the $u_{i}$ are represented statistically and responsible for calculating TTPs’ confidence $C$, allowing quantify trust based on the diversity of trust attributes $j$.
$pos_{ki}^{(\tau)}$ $=\mathcal{A}\cdot\sum_{j=1}^{J}pos_{ki}^{j},\quad neg_{ki}^{(\tau)}$ $=\mathcal{A}\cdot\sum_{j=1}^{J}neg_{ki}^{j}$ (10)
To enhance the precision of area relevance ($AR$) assessments within our trust model, we employ Prompt Engineering to direct the inference process of multiple LLMs debating $\mathcal{D}$. The robustness of our final decision is strengthened by integrating semantic similarity measures, denoted as $ST$. Specifically, we employ the all-mpnet-base-v2 sentence transformer in our cases, which yields semantically rich embeddings that effectively capture nuanced linguistic relations [32]. This method refines the evaluation of area relevance between a trustee’s professional area, as verified by TTPs, and the sensitivity required by the prompts. The formula for computing $AR$ in $[0,1]$ is presented as follows:
@@ 4.2 Authority-Verified Trust
Remark 1: We also incorporate periodic re-validation throughout the whole trust evaluation process, which is a common practice in web services [2], to ensure that all sensitive information requests meet specific security requirements within guardrail systems. This mandatory procedure applies universally, irrespective of trust scores, preventing outdated or overly trusted credentials from granting inappropriate access and thus enhancing the system’s robustness. For long-term access to sensitive or controversial information for legal or specialized purposes, users must periodically re-validate their identity with TTPs to maintain credential validity.
@@ 4.3 Contextual Adaptation Mechanism
Guardrails % In-domain Questions % Out-of-domain Questions
@@ 5.1 Data Preparation
For our experiments, we gathered data from several datasets including AdvBench [45], HarmBench [26] and ToxicChat [22], categorizing prompts by their relevance to different domains to evaluate our adaptive trust model for guardrails. This dataset comprises 520 harmful prompts reflecting potential sensitive interactions, supplemented by 3,000 safe prompts for interaction diversity. We partitioned the prompts into in-domain and out-of-domain categories based on whether their topics matched the user’s professional domain. For example, prompts related to computer science were classified as in-domain for users from the computer science (CS) domain, while prompts from other fields were considered out-of-domain. Due to privacy concerns and restricted access to actual user data from TTPs, we created synthetic profiles for virtual users, incorporating simulated ratings, professional areas, social information, and more, to emulate real-world trust attributes, allowing us to verify if the trust model effectively aligns access to sensitive information with validated user identities. All necessary materials including code and datasets, are hosted on GitHub [^8] The anonymized GitHub will be made public upon paper acceptance to comply with double-blind review requirements.
5.2 Implementation Configurations